OpenStudy (anonymous):

What is the solution to the system of equations? 4x+y=2 x-y=3 ( , )

OpenStudy (anonymous):

Do you have any idea of where to start?

OpenStudy (usukidoll):

elimination method.....

OpenStudy (anonymous):

substitution method

OpenStudy (anonymous):

The answer I got is (1, -2), if you would like more help on this I will help you.

OpenStudy (anonymous):

thank you can you help me with more???

OpenStudy (anonymous):

What is the solution to the system of equations? y=2x+6 3x-y=6

OpenStudy (usukidoll):

substitution method

OpenStudy (anonymous):

yes

OpenStudy (usukidoll):

the solve for x

OpenStudy (anonymous):

i should find x and y =p

OpenStudy (usukidoll):

ya

OpenStudy (anonymous):

You get the answer fine for that one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Any more? For the last one I got (12, 30)

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

Solve the given system, using the substitution method. y = 3x – 7 6x – 2y = 12 A. (12, 14) B. There are an infinite number of solutions. C. There is no solution. D. (14, 12)

OpenStudy (anonymous):

y=3x-7 6x-2y=12

OpenStudy (anonymous):

Well looking at this the only one I think makes sense is that its going to be no solution exists.

OpenStudy (anonymous):

Solve, using the substitution method. y – 3x = 7 21 + 9x = 3y A. There is no solution. B. There are an infinite number of solutions. C. The solution is (1, 10) D. The solution is (28, 0)

OpenStudy (usukidoll):

*farts*

OpenStudy (anonymous):

y-3x=7 21+9x=3y

OpenStudy (anonymous):

>.< Lol

OpenStudy (anonymous):

I'm not one hundred percent sure on this I'm leaning toward infinite number of solutions, any idea @usukiDoll

OpenStudy (anonymous):

One more ??!!

OpenStudy (anonymous):

Solve, using the substitution method. y + 2x = 7 14 – 4x = 2y A. The solution is (1, 5) B. The solution is (21, 0) C. There are an infinite number of solutions. D. There is no solution.

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

y + 2x = 7 14 – 4x = 2y

OpenStudy (usukidoll):

omg solve this with subsitution method . we know that the system is inconsistent if we have something that's false and all of the variables vanish like \[0 \neq 2\]

OpenStudy (usukidoll):

OK LET'S GO ONE AT A TIME!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Could you clarify on the set since there's a bunch of ???'s

OpenStudy (anonymous):

ok y+2x=7 14-4x=2y

OpenStudy (usukidoll):

 it is many solutions.. we got lucky